OpenStudy (anonymous):

g(t) = 1/t +1 - find g^-1(t)

OpenStudy (anonymous):

i got y = t-1

OpenStudy (anonymous):

but that cant be right

hartnn (hartnn):

u should get 1/(t-1)

OpenStudy (anonymous):

\[t=\frac{ 1 }{ y }+1\]

OpenStudy (anonymous):

I don't think you want that "y" in there...

hartnn (hartnn):

y=1/t+1 y-1 =1/t t=1/(y-1) now replace

OpenStudy (anonymous):

but cant u just replace y for t at the beginning to do inverses? or is that harder?

hartnn (hartnn):

same thing

OpenStudy (anonymous):

yeah, you're right... as you wrote it 2 up, looks good to me

OpenStudy (anonymous):

the final form shouldn't have the y, though, should it? Should just be function of t, right?

OpenStudy (anonymous):

right.. the answer would be \[f ^{-1}(x) = \frac{ 1 }{ t-1 }\] i am having serious issues with this for whatever reason.

hartnn (hartnn):

what issues ? didn't u get how it came ?

OpenStudy (anonymous):

embarrasingly enough i cannot figure out the fraction.. lol

hartnn (hartnn):

t=1/y+1 t-1 =1/y y=1/(t-1)

OpenStudy (anonymous):

y = 1/t +1 . so i subtract 1 ofrom each side.. y-1 = 1/t .. now? multiply y to each side? thats where i get confused.

OpenStudy (anonymous):

i mean multiply t to each side to get rid of the fraction?

hartnn (hartnn):

take reciprocal

hartnn (hartnn):

if a = 1/b, then 1/a = b

OpenStudy (anonymous):

did u skip a step between t-1 = 1/y and your answer?

hartnn (hartnn):

just taking reciprocals.........

OpenStudy (anonymous):

meaning.. from t-1 = 1/y... i should multiply both sides by the reciprocal. y/1?

hartnn (hartnn):

if a = 1/b, then 1/a = b so if t-1 = 1/y then y=1/(t-1)

OpenStudy (anonymous):

so u dont multiply both sides by the reciprocal?

hartnn (hartnn):

u want to multiply or divide something ?? 1. multiply both sides by y 2. then divide both sides by t-1

OpenStudy (anonymous):

sorry i just dont have that immeidate association with math theorems

OpenStudy (anonymous):

so a=1/b then 1/a = b makes sense.. but ...

OpenStudy (anonymous):

y(t-1) = 1 then divide both sides by t-1. awesome thanks.

hartnn (hartnn):

i m glad that u got it :)